OpenStudy (anonymous):

Which property justifies the following equation? 7[6+5+(-6)]=7[6+(-6)+5] A. distributive B. commutative C. associative D. identity

OpenStudy (anonymous):

i would like to say associative property but give me a moment

OpenStudy (anonymous):

Its not Distributive or identity.. Do you have a clue which one it is (Commutative or Associative) @Angel06061x

OpenStudy (anonymous):

ok so it wouldnt be identity nor distrubitive

OpenStudy (anonymous):

Its Associative... Do you know why?????

OpenStudy (anonymous):

a+b=b+a (commutative property )

OpenStudy (anonymous):

because its the same thing but just flipped?

OpenStudy (anonymous):

if u see the images of each property u will for sure know why with commutative think of it as a mirror and associative think of parenthesis

OpenStudy (anonymous):

What they said^^^

OpenStudy (anonymous):

okay thank yous

OpenStudy (anonymous):

5+(-6)=(-6)+5